Section 4: Using Graphs Use the graph belox to answer questions 2-23 10 Price Floor Doriginal 8 10 26. Use the information above to find the demand equation. Suppose a Price floor is implanted in the market above at 56.00. 27. Is the price floor binding? Explain why 28. Calculate the deadweight loss caused by the price floor. 29. Calculate the transfer of surplus from the consumer to the producer.
Answer #1

10 10 price floor 10 Transfer of CS 14 16 18 20 PE-0.5Q 10 0 2 46 8 10 12 14 16 18 20 22 Quantity 7

26. From the graph we get the inverse relation as below
P = -0.5Q+10
0.5Q=10-P
Q= 20-2P

27.Yes, as the price floor is set above the equilibrium price, it is binding.

28. DWL = area of the triangle = 0.5*(6-4)*(10-8)= 0.5*2*2 = 0.5*4=2

29. area of the rectangle = 8*(6-4) =8*2=16